# Transfer tensor analysis of localization in the Anderson and Aubry-André-Harper models

### v1

@@ IV Summary
Applying the transfer tensor methods to simple models of transport in lattices has demonstrated how memory effects differentiate static and dynamic disorder while employing much smaller bases than are needed for conventional numerical studies of transport properties. The structure of the tensors explains the meaning of the infinite memory which arises when performing an ensemble average over static disorder and restoring spatial homogeneity. Memory removes fictitious terms that would correspond to redrawing static disorder thereby converting it to dynamic disorder. This eternal memory, while necessary for localization, is not a sure indication of it. However, localization can be inferred from the total outgoing-pseudoflux. The outgoing-pseudoflux behaves similarly to the diffusion constant in the Anderson model, perhaps indicating that more complicated metrics of the tensors might produce the diffusion constant itself or indicate whether a system is diffusive or ballistic. Future inquiries in this direction are warranted as transfer tensor methods show promise as an analysis tool for understanding memory effects and addressing transport properties.
V Acknowledgments
@@ VI Data Availability
VII Appendix A: Steady-state analysis of transfer tensors for localized systems
@@
For a finite, delocalized system subject to the HSR infinite temperature bath, we expect the steady-state solution to be $\rho^{(ss)}_{nn}=p_{n}^{(ss)}=1/N$, where $p_{n}^{(ss)}$ is the steady state population of site $n$, regardless of the initial state.[61] On the other hand, for localized systems, we anticipate an initial state-dependent steady state solution that is generally spatially inhomogeneous, $\rho_{nn}^{(ss)}\neq\rho_{mm}^{(ss)}$. By substituting $\rho^{(ss)}$ into Eq. (3), we have
@@
Figure 3: Systems with exponentially correlated noise ($\langle\epsilon_{n}(t)\epsilon_{m}(t^{\prime})\rangle=\Gamma e^{-\gamma|t-t^{\prime}|}\delta_{nm}$) with 9 sites, $V\Delta t=1/10$, $\Gamma/V=1$, and various $\gamma$ values. (a) Population at the central site as a function of time step and $\gamma$. The full thermalization limit $\rho_{m,m}(t\rightarrow\infty)=N^{-1}$ is indicated by the dashed line. (b) Tensor Frobenius norm. The plateauing for smaller $\gamma$ curves suggest more noise samples are needed for convergence. With more noise samples the plateau continues to decrease in value. (c) Outgoing-pseudoflux.
VIII Appendix B: Correlated Noise
@@
The interplay between static disorder and dynamic noise manifested in the mixed Anderson/HSR model can be analogously present in systems with correlated noise. Disorder can be thought of as perfectly correlated, constant noise, while the HSR white noise is the opposite limit with vanishing temporal correlation. Here, we demonstrate this analogy using a system with temporally exponentially correlated (but spatially independent) on-site energies, i.e. $\langle\epsilon_{n}(t)\epsilon_{m}(0)\rangle=\Gamma e^{-\gamma t}\delta_{nm}$. With this model, two free parameters $\Gamma$, the noise strength, and $\gamma$, the inverse correlation time exist. In the limits of $\gamma\rightarrow\infty$ and $\gamma\rightarrow 0$ one recovers the HSR and the Anderson models, respectively.
@@
The numerical results are shown in Fig. 3 in the same convention as in Fig. 1. As predicted, the dynamics and the corresponding tensors follow the same trend as in the mixed Anderson/HSR models studied in the main text. In Fig.  3 (a) we examine the convergence of the dynamics and the TT in the short time limit for systems with same noise strength $\Gamma$ and varying correlation time $\gamma^{-1}$. In part (b) the tensor norms, after a certain transient regime around $k=30$ for larger $\gamma$ values, decay single exponentially in accordance with the rates $\gamma$. Finally, we show the outgoing-pseudoflux in part (c). In contrast to the mixed Anderson/HSR models with constant disorder and increasing noise strength shown in Fig.  1 (d), here the pseudoflux increases monotonically with increasing $\gamma$.

### v2

@@ IV Summary
Applying the transfer tensor methods to simple models of transport in lattices has demonstrated how memory effects differentiate static and dynamic disorder while employing much smaller bases than are needed for conventional numerical studies of transport properties. The structure of the tensors explains the meaning of the eternal memory which arises when performing an ensemble average over static disorder and restoring spatial homogeneity. Memory removes fictitious terms that would correspond to redrawing static disorder thereby converting it to dynamic disorder. This eternal memory, while necessary for localization, is not a sure indication of it. However, localization can be inferred from the total outgoing-pseudoflux. The outgoing-pseudoflux behaves similarly to the diffusion constant in the Anderson-HSR model, perhaps indicating that more complicated metrics of the tensors might produce the diffusion constant itself or indicate whether a system is diffusive or ballistic. Future inquiries in this direction are warranted as transfer tensor methods show promise as an analysis tool for understanding memory effects and addressing transport properties.
Acknowledgment
@@
Author Contributions
Michelle C. Anderson and Chern Chuang contributed equally to this work.
Appendix A: Steady-state analysis of transfer tensors for localized systems
@@
For a finite, delocalized system subject to the HSR infinite temperature bath, we expect the steady-state solution to be $\rho^{(ss)}_{nn}=p_{n}^{(ss)}=1/N$, where $p_{n}^{(ss)}$ is the steady state population of site $n$, regardless of the initial state.[note2] On the other hand, for localized systems, we anticipate an initial state-dependent steady state solution that is generally spatially inhomogeneous, $\rho_{nn}^{(ss)}\neq\rho_{mm}^{(ss)}$. By substituting $\rho^{(ss)}$ into Eq. (3), we have
@@
The current framework provides limited insight into the numerical correspondence of the outgoing-pseudoflux and the diffusion constant observed in the Anderson-HSR model, while the correlation between the two are strongly implicated by their similar non-monotonic behavior with respect to $\Gamma$ as shown in Fig. 2. Some intuition can be developed from the Green-Kubo expression for the diffusion constant
$D=\frac{1}{Z}\int_{0}^{\infty}dt~\mathrm{Tr}[e^{-\beta H}\hat{j}(t)\hat{j}(0)]$ (12)
where $\hat{j}(t)$ is the flux operator and $e^{-\beta H}/Z=N^{-1}$ is the partition function for the infinite-temperature HSR model.[25] The time-evolution of operators in the Heisenberg picture can be similarly defined as in Eq. (3) albeit with a complex conjugate. In the current model of nearest-neighbor tight-binding chain, we have $\hat{j}=iV\sum_{n}|n\rangle\langle n+1|+H.c.$. With this, it becomes clear that the integrand in Eq. (12) is in fact related to a quantity different from $F_{m}(k\rightarrow\infty)$:
$\frac{1}{N}\mathrm{Tr}[\hat{j}(k\Delta t)\hat{j}(0)]=2V^{2}\sum_{n}\left[M_{n+1,n,1,0}(k)-M_{n-1,n,1,0}(k)\right].$ (13)
And the diffusion constant can be approximated by the $k$-summation of this quantity in the $\Delta t\ll 1$ limit. Note that disorder-averaged dynamical maps $M(k)$’s are used, instead of TT. This is indicative of the physical significance and the distinction between the maps and TT. To track the temporal evolution of a dynamical quantity (flux in this case), one needs to account for all possible pathways leading to the current time step, Markovian or not, signified by the nested tensor multiplications in Eq. (3).
@@
Figure 3: Systems with exponentially correlated noise ($\langle\epsilon_{n}(t)\epsilon_{m}(t^{\prime})\rangle=\Gamma e^{-\gamma|t-t^{\prime}|}\delta_{nm}$) with 9 sites, $V\Delta t=0.1$, $\Gamma/V=1$, and various $\gamma$ values. (a) Population at the central site as a function of time step and $\gamma$. The full thermalization limit $\rho_{m,m}(t\rightarrow\infty)=N^{-1}$ is indicated by the dashed line. (b) Tensor Frobenius norm. The plateauing for smaller $\gamma$ curves suggest more noise samples are needed for convergence. With more noise samples the plateau continues to decrease in value. (c) Outgoing-pseudoflux.
Appendix B: Correlated Noise
@@
The interplay between static disorder and dynamic noise manifested in the mixed Anderson-HSR model can be analogously present in systems with correlated noise. Disorder can be thought of as perfectly correlated, constant noise, while the HSR white noise is the opposite limit with vanishing temporal correlation. Here, we demonstrate this analogy using a system with temporally exponentially correlated (but spatially independent) on-site energies, i.e. $\langle\epsilon_{n}(t)\epsilon_{m}(0)\rangle=\Gamma e^{-\gamma t}\delta_{nm}$. With this model, two free parameters $\Gamma$, the noise strength, and $\gamma$, the inverse correlation time exist. In the limits of $\gamma\rightarrow\infty$ and $\gamma\rightarrow 0$ one recovers the HSR and the Anderson models, respectively.
@@
The numerical results are shown in Fig. 3 in the same convention as in Fig. 1. As predicted, the dynamics and the corresponding tensors follow the same trend as in the mixed Anderson-HSR models studied in the main text. In Fig.  3 (a) we examine the convergence of the dynamics and the TT in the short time limit for systems with same noise strength $\Gamma$ and varying correlation time $\gamma^{-1}$. In part (b) the tensor norms, after a certain transient regime around $k=30$ for larger $\gamma$ values, decay single exponentially in accordance with the rates $\gamma$. Finally, we show the outgoing-pseudoflux in part (c). In contrast to the mixed Anderson-HSR models with constant disorder and increasing noise strength shown in Fig.  1 (d), here the pseudoflux increases monotonically with increasing $\gamma$.